You are considering buying a $8,000 bond with a bond rate 6% paying once per year and a 15 year maturation period. You would like 10% annual yield and you plan to keep until maturity. How much should you pay for the bond?
Answer #1

Price to be paid = Present value of future inflows @ 10%

Bond interest per year = Bond face value * Bond rate = $8,000 * 6% = $480.

Price to be paid = Present value of face value of bond + Present value of future interest cash inflows

Present value of face value of bond = Bond's face value * Present value factor @10% for 15th year = $8,000 * 0.2394 = $1,915.2

Present value factor @10% for 15th year = 1 / [ (1 + 0.10)^15 ]

Present value of future interest cash inflows = Interest amount * Present value annuity factor @10 % for 15 years = $480 * 7.6061 = $3,650.9

Price to be paid = $1,915.2 + $3,650.9 = $5,566.1

For present value factor - Please see the present value factor and present value annuity factor table